The first of these equations plus the second equals the third: x+ y+ z=2 X + 2y+ Z = 3 2x + 3y + 2z = 5. The first two planes meet along a line. The third plane contains that line, because if x, y, z satisfy the first two equations then they also __ . The equations have infinitely many solutions (the whole line L). Find three solutions on L.  


1
Expert's answer
2021-04-27T13:41:39-0400

We have given three equations,

"x+y+z =2 -----(1)"

"x+2y+z =3 ----(2)"

"2x+3y+2z = 5 -----(3)"

We can observe that,

"(1)+(2) = (3)"

If "x,y,z" satisfy "(1),(2)" they satisfy "(3)" also.

So, (3) also contains the line of intersection of (1) and (2) planes.

"(2)- (1)"

This gives "y=1" "\implies x+1+z = 2 \implies x+z = 1"

Let "x=0 \implies z=1"

Let "x=1 \implies z=0"

Let "x=-1 \implies z=2"

Solutions are "(0,1,1),(1,1,0){}and {}(-1,1,2)"